Enfield has its vehicle assembly plant in Tiruvottiyur, Chennai and its

engine assembly plant in Oragadam near Chennai. Engines are transported between the

two plants using trucks, with each trip costing Rs. 2000. vehicle plant assembles and

sells 250 motorcycles every day. Each engine costs Rs. 25000, and MCL incurs a holding

cost of 24 % per year.

 

a. How many engines should MCL load onto each truck? what is the optimal order

quantity?

 

b. What is the cycle inventory of engines at MCL? If demand, and thus production, for

their motorcycle grows, but all other input data remain unchanged, would you expect

cycle inventory of engines at MCL to increase or decrease?



1
Expert's answer
2021-05-04T12:16:53-0400

a. Optimal order quantity is:

"EOQ = \sqrt{\frac{2×2,000×250}{25,000×0.24}} = 12.91."

b. An inventory cycle, also called inventory turnover, or turns, is a measure of how many times during a given period a company sells an volume of products equal to the volume of supplies it keeps on hand.

Inventory cycle = 1/0.24 = 4.17 units.

If demand, and thus production, for their motorcycle grows, but all other input data remain unchanged, then we would expect cycle inventory of engines at MCL to decrease.